### v1

@@ III-A Target Graph Construction
[…]
@@
      11 if  ${\tt distance}(n_{i},{\boldsymbol{N}})>a$ then  ${\boldsymbol{N}}\leftarrow{\tt add\_node}(n_{i})$
@@
We need to take special consideration for the different formed angles by cracks. Fig. 4 illustrates a few examples of $\mathbb{G}_{c}$ with different formed angles. Fig. 4 shows a general case where two vertex points (i.e., $N_{1}$ and $N_{2}$) are far away with a distance larger than $a$. When a vertex point on $\mathcal{T}$ forms an acute angle, it can overlap with its own Minkowski sum area, which makes the nodes visible to each other; see Fig. 4. In this case, the shortest path does not transverse the crack, which leads to unfilled cracks. To achieve the shortest crack-filling path, the vertex must be at least a distance of $a$ from the boundary of the Minkowski sum area. These cracks are pre-identified, and their respective Minkowski offset values are adjusted (line 5 in Algorithm 1). Fig. 4 shows the Minkowski sum area with the adjusted Minkowski offset value. By doing so, it guarantees that the calculated shortest path covers the entire crack in ${\tt shortest\_path}$.
@@ III-B Target Coverage Planning
With $\mathbb{G}_{c}$, we consider the required properties of the graph to guide the robot to traverse the cracks with minimum cost. First, all of the graph vertices except the first and last in the route must have an even number of connected edges; otherwise, no route exists in the graph to allow traveling along each edge exactly once; that is, the robot would get stuck at odd vertices. To prevent this, the vertices with an odd number of connected edges are made even by adding edges. To achieve the shortest traveling distance, the robot motion planner must search for all the different means to pair off the vertices with odd numbers of connected edges and choose the pair that adds the least total distance to the graph.
@@
The task of finding the shortest route covering all the graph edges is similar to the Chinese Postman Problem (CPP) [37]. However, the crack edges might not form a connected graph and instead form subsets of graphs. As a result, when such situations arise, we adapt the rural postman problem (RPP) solution [38]. The RPP was demonstrated to be NP-complete, and heuristic solution procedures were proposed to approximate the solution [39]. Unlike any postman problem in which the postman usually needs to finish at the same location as the starting point, robot $\mathbf{R}$, however, can start and finish the job at different locations. Therefore, two nodes that have an odd number of connected edges are left unpaired and selected as the starting and ending positions, respectively.
Algorithm 2 illustrates the ${\tt GCC}$ planning algorithm. The algorithm formulates an RPP problem to guarantee that the robot covers all the cracks with the least number of revisits. The algorithm comprises three phases. In the first phase, the minimum spanning tree is computed over $\mathbb{G}_{c}$ to get a subset of the edges $\mathbb{E}_{\text{t}}$ that connects all the vertices with the minimum possible total edge weight (line 1). The union of $\mathbb{G}_{c}$ and $\mathbb{E}_{\text{t}}$ ensures a single connected network for the postman problem. In the second phase, using matching theory [40], we search for all possible ways to pair up the odd vertices, $N_{\text{odd}}$ (line 2) by describing the solution as a linear programming (LP) polyhedron. ${\boldsymbol{E}}_{\text{add}}$ represents the collection of all sets of added edges for every potential pairing, while $\mathbb{E}_{\text{add}}(i)$ indicates the added edge for the $i$th potential pairing. $C_{\text{total}}$, $C_{\text{max}}$, and $C_{\text{final}}$ denote the total, maximum, and final edge cost, respectively. The edge cost is formulated as the line distance. $E_{\text{final}}$ is selected as the minimum cost of the total added edges minus the maximum-cost single added edge. The maximum-cost single edge is broken into the starting and ending nodes of the path (lines 3 to 6). The graph is updated by adding $E_{\text{final}}$ (line 7). In the final phase, the function ${\tt fleury}$ [41] is used to obtain the optimal path $\mathcal{P}_{c}$ for $\mathbf{R}$. For example, as shown in Fig. 3, the optimal path is obtained as a sequence $\mathcal{P}_{c}:N_{4}\rightarrow N_{1}\rightarrow N_{2}\rightarrow N_{6}\rightarrow N_{5}\rightarrow N_{3}\rightarrow N_{5}\rightarrow N_{7}\rightarrow N_{9}\rightarrow N_{10}\rightarrow N_{11}\rightarrow N_{9}\rightarrow N_{7}\rightarrow N_{6}\rightarrow N_{12}\rightarrow N_{8}$.
@@
21 $\mathbb{E}_{\text{t}}\leftarrow{\tt MST}(\mathbb{G}_{c})$, $\mathbb{G}_{c}\leftarrow\mathbb{G}_{c}\cup\mathbb{E}_{\text{t}}$ $N_{\text{odd}}\leftarrow{\tt find\_oddNode}(\mathbb{G}_{c})$, ${\boldsymbol{E}}_{\text{add}}\leftarrow{\tt pair}(N_{\text{odd}},\mathbb{G}_{c})$ for each ${\boldsymbol{E}}_{\text{add}}\rightarrow{\mathbb{E}_{\text{add}}}(i)$ do
@@
[…]
@@ IV-A ${\tt SCC}$ Planner
To explain our approach, we borrow the following definitions from [19, 21]. As shown in Fig. 5, the slice is effectively a vertical line sweeping from left to right along the sweep direction in $\mathcal{W}$. A cell is an area where slice connectivity does not change, and changes in the connectivity of the slice only occur at critical points. A critical point is located on the boundary of an object whose surface normal is perpendicular to the sweep direction. Critical points are used to determine the cell boundaries. A target region, denoted by $\mathcal{M}_{\text{c}}$, is obtained by a dilated $\mathbb{G}_{c}$ by a circular disk area with a radius of $S$. For example, $\mathcal{M}_{\text{c}}$ is the yellow area in Fig. 5.
Unlike most MCD-based coverage path planning problems, the critical points defined here are not only on the boundary of obstacles but also on the boundary of the target regions. According to the MCD of free space $\mathcal{W}$, a Reeb graph [21, 19] is constructed and denoted as $\mathbb{G}_{\text{w}}$. In Fig. 5, $\mathbb{G}_{\text{w}}=({\boldsymbol{C}},{\boldsymbol{E}})$ is shown as the black solid lines. The nodes of $\mathbb{G}_{\text{w}}$ are the critical points ${\boldsymbol{C}}=\{C_{i}\}$, and the edges ${\boldsymbol{E}}=\{E_{i}\}$ connect the neighboring critical points. The edge $E_{i}$ in the Reeb graph directly corresponds to the cell $\mathcal{A}_{i}$ in the free space, $i=1,\cdots,6$, in the figure.
@@
Output : Path $\mathcal{P}_{R}$
[…]
@@
Algorithm 3 briefly describes the ${\tt SCC}$ planner. With $\mathbb{G}_{\text{w}}$ and $\mathbb{G}_{c}$, we search the shortest cost route covering all the graph edges at least once; that is, we connect edges in both graphs to form an Euler tour with the least cost. In the case of an unspecified ending position, all of the graph nodes except for the initial and ending vertices must have an even number of connected edges (i.e., an even degree); otherwise, the robot could get stuck at vertices with odd degrees. Therefore, the degree of vertices is maintained even by adding edges. For example, as shown in Fig. 5, $C_{1}$-$C_{6}$, $N_{1}$, and $N_{3}$-$N_{5}$ are the vertices with odd degrees, and the red dashed edges connecting {$C_{2}$,$N_{1}$}, {$C_{3}$,$N_{3}$}, {$C_{5}$,$N_{4}$}, and {$C_{4}$,$N_{5}$} are added to the graph to form the shortest route. The resulting Euler tour, denoted as $\mathcal{P}_{\text{wc}}$, provides an order of edges that the robot should visit (line 2). The connecting distance between each adjacent cell in the Euler tour is minimized in the function ${\tt cell\_connect}$. The variable conn represents the connections of adjacent cells in the Euler tour (line 3). Finally, complete coverage is performed by following the sequence of edges in the Euler tour (line 4).
@@ IV-B ${\tt oSCC}$ Planner
When target information is unknown, robot $\mathbf{R}$ needs to detect all targets and the critical points online in $\mathcal{W}$. We first combine graphs $\mathbb{G}_{\text{c}}$ and $\mathbb{G}_{\text{w}}$ using the following lemma and propositions with proofs given in Appendices A and B.
@@
[…]
@@
[…]
@@
Output : Path $\mathcal{P}_{R}$
1 $\mathcal{M}_{\text{cov}}\leftarrow\O$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$ while $\mathcal{M}_{\text{cov}}\neq\mathcal{W}$ do

### v2

@@ III-A Target Graph Construction
Algorithm 1 describes the construction of graph $\mathbb{G}_{c}$. The input is the binary image $\mathcal{I}$, which is homography warped according to their actual shape. The first step is to use the skeleton method to find the topology $\mathcal{T}$ of $\mathcal{I}$ (line 1). Endpoints of the cracks, $P_{\text{end}}$, are found from $\mathcal{T}$ by searching the 8 neighbors of each pixel. $\mathcal{T}$ is then separated into individual cracks, $\textit{crack}_{\mathcal{I}}$, by the endpoints and crack extension direction (line 2). Lines 3 and 4 find all $P_{\text{end}}$, with a distance larger than $a$ between each other to form end-nodes $N_{\text{end}}$. Footprint region $\mathcal{M}_{\text{f}}$ is obtained by the Minkowski sum (line 5). The intersected areas $A_{\text{int}}$ of all dilations are found, and their corresponding overlapping numbers $\textit{No}_{\text{over}}$, and centroid point $P_{\text{int}}$ are determined (lines 6-7). Among $P_{\text{int}}$ and $N_{\text{end}}$ being the final candidates for the graph nodes $N_{\text{cand}}$ (line 9), we select the nodes with a distance greater than $a$ with each other as the nodes ${\boldsymbol{N}}$ of $\mathbb{G}_{c}$ (line 11). The graph’s edges ${\boldsymbol{E}}_{c}$ are obtained from its nodes and crack topology (line 12). Function ${\tt shortest\_path}$ is used to adjust the graph edges to guarantee their locations inside the Minkowski area by using visibility graph shortest path planning (line 13).
@@
1 $\mathcal{T}\leftarrow{\tt topology}(\mathcal{I})$, $P_{\text{end}}\leftarrow{\tt get\_endpoint}(\mathcal{T})$
2 $\textit{crack}_{\mathcal{I}}\leftarrow{\tt get\_cracks}(\mathcal{T},P_{\text{end}})$
for each $\text{each }P_{\text{end}}\rightarrow e_{i}$ do
      3 $P_{\text{test}}\leftarrow P_{\text{end}}/e_{i}$
       if  $\text{all}({\tt distance}(e_{i},P_{\text{test}}))>a$ then
@@ 3 $P_{\text{test}}\leftarrow P_{\text{end}}/e_{i}$
6 $(A_{\text{int}},\textit{No}_{\text{over}})\leftarrow{\tt get\_intersection}(\mathcal{M}_{\text{f}})$
7 $P_{\text{int}}\leftarrow{\tt centroid}(A_{\text{int}})$
8 $(\textit{crack}_{\mathcal{I}},N_{\text{end}})\leftarrow{\tt shorten}(\textit{crack}_{\mathcal{I}},N_{\text{end}})$
9 $N_{\text{cand}}\leftarrow{\tt combine}(P_{\text{int}},N_{\text{end}})$
10 $N_{\text{cand}}\leftarrow{\tt sort}(N_{\text{cand}},\textit{No}_{\text{over}})$
for each $\text{each }N_{\text{cand}}\rightarrow n_{i}$  do
       if  ${\tt distance}(n_{i},{\boldsymbol{N}})>a$ then
@@ 10 $N_{\text{cand}}\leftarrow{\tt sort}(N_{\text{cand}},\textit{No}_{\text{over}})$
13 $\mathbb{G}\leftarrow\{{\boldsymbol{N}},{\boldsymbol{E}}_{c}\}$, $\mathbb{G}_{c}\leftarrow{\tt shortest\_path}(\mathbb{G},\mathcal{M}_{\text{f}})$
@@
We need to take special consideration for the different formed angles by cracks. Fig. 4 illustrates a few examples of $\mathbb{G}_{c}$ with different formed angles. Fig. 4 shows a general case where two vertex points (i.e., $N_{1}$ and $N_{2}$) are far away with a distance larger than $a$. When a vertex point on $\mathcal{T}$ forms an acute angle, it overlaps with its own Minkowski sum area, which makes the nodes visible to each other; see Fig. 4. In this case, the shortest path does not transverse the crack, which leads to unfilled cracks. To achieve the shortest crack-filling path, the vertex must be at least a distance of $a$ from the boundary of the Minkowski sum area. These cracks are pre-identified, and their respective Minkowski offset values are adjusted (line 5 in Algorithm 1). Fig. 4 shows the Minkowski sum area with the adjusted Minkowski offset value. By doing so, it guarantees that the calculated shortest path covers the entire crack in ${\tt shortest\_path}$.
@@ III-B Target Coverage Planning
With $\mathbb{G}_{c}$, we consider the required properties of the graph to guide the robot to traverse the cracks with minimum cost. All the graph vertices except the first and last in the route must have an even number of connected edges. Otherwise, no route exists in the graph to allow traveling along each edge exactly once, meaning the robot would get stuck at odd vertices. To prevent this, vertices with an odd number of connected edges are made even by adding edges. To achieve the shortest traveling distance, the robot motion planner must search for all the different ways to pair off the vertices with odd numbers of connected edges and choose the pair that adds the least total distance to the graph.
@@
The task of finding the shortest route covering all the graph edges is similar to the Chinese Postman Problem (CPP) [37]. However, the crack edges might not form a connected graph and instead form subsets of graphs. As a result, when such situations arise, we adapt the rural postman problem (RPP) solution [38]. The RPP was demonstrated to be NP-complete, and heuristic solution procedures were proposed to approximate the solution [39]. Unlike in any postman problem where the postman typically needs to finish at the same location as the starting point, robot $\mathbf{R}$ can start and finish the job at different locations. Therefore, two nodes with an odd number of connected edges are left unpaired and selected as the starting and ending positions, respectively.
Algorithm 2 illustrates the ${\tt GCC}$ planning algorithm. The algorithm formulates an RPP problem to guarantee that the robot covers all the cracks with the least number of revisits. The algorithm comprises three phases. In the first phase, the minimum spanning tree is computed over $\mathbb{G}_{c}$ to get a subset of the edges $\mathbb{E}_{\text{t}}$ that connects all the vertices with the minimum possible total edge weight (line 1). The union of $\mathbb{G}_{c}$ and $\mathbb{E}_{\text{t}}$ ensures a single connected network for the postman problem. In the second phase, using matching theory [40], we search for all possible ways to pair up the odd vertices $N_{\text{odd}}$ (line 2) by describing the solution as a linear programming (LP) polyhedron. ${\boldsymbol{E}}_{\text{add}}$ represents the collection of all sets of added edges for every potential pairing, while $\mathbb{E}_{\text{add}}(i)$ indicates the added edge for the $i$th potential pairing. $C_{\text{total}}$, $C_{\text{max}}$, and $C_{\text{final}}$ denote the total, maximum, and final edge cost, respectively. The edge cost is formulated as the line distance. $E_{\text{final}}$ is selected as the minimum cost of the total added edges minus the maximum-cost single added edge. The maximum-cost single edge is broken into the starting and ending nodes of the path (lines 3 to 6). The graph is updated by adding $E_{\text{final}}$ (line 7). In the final phase, the function ${\tt fleury}$ [41] is used to obtain the optimal path $\mathcal{P}_{c}$ for $\mathbf{R}$. For example, as shown in Fig. 3, the optimal path is obtained as a sequence $\mathcal{P}_{c}:N_{4}\rightarrow N_{1}\rightarrow N_{2}\rightarrow N_{6}\rightarrow N_{5}\rightarrow N_{3}\rightarrow N_{5}\rightarrow N_{7}\rightarrow N_{9}\rightarrow N_{10}\rightarrow N_{11}\rightarrow N_{9}\rightarrow N_{7}\rightarrow N_{6}\rightarrow N_{12}\rightarrow N_{8}$.
@@
1 $\mathbb{E}_{\text{t}}\leftarrow{\tt MST}(\mathbb{G}_{c})$, $\mathbb{G}_{c}\leftarrow\mathbb{G}_{c}\cup\mathbb{E}_{\text{t}}$
2 $N_{\text{odd}}\leftarrow{\tt find\_oddNode}(\mathbb{G}_{c})$, ${\boldsymbol{E}}_{\text{add}}\leftarrow{\tt pair}(N_{\text{odd}},\mathbb{G}_{c})$
for each ${\boldsymbol{E}}_{\text{add}}\rightarrow{\mathbb{E}_{\text{add}}}(i)$ do
      3 $C_{\text{max}}\leftarrow{\tt max}({\mathbb{E}_{\text{add}}}(i).\textit{edgeCost})$
      4 $C_{\text{total}}(i)\leftarrow{\tt sum}({\mathbb{E}_{\text{add}}}(i).\textit{edgeCost})-C_{\text{max}}$
      5 $(C_{\text{final}},\textit{Index})\leftarrow{\tt min}(C_{\text{total}})$
@@
7 $\mathbb{G}_{c}\leftarrow{\tt add}(\mathbb{G}_{c},E_{\text{final}})$, $\mathcal{P}_{c}\leftarrow{\tt fleury}(\mathbb{G}_{c})$
@@ IV-A ${\tt SCC}$ Planner
To explain our approach, we borrow the following definitions from [19, 21]. As shown in Fig. 5, the slice is effectively a vertical line sweeping from left to right along the sweep direction in $\mathcal{W}$. A cell is an area where slice connectivity does not change, and changes in the connectivity of the slice only occur at critical points. A critical point is located on the boundary of an object whose surface normal is perpendicular to the sweep direction. Critical points are used to determine the cell boundaries. A target region, denoted by $\mathcal{M}_{\text{c}}$, is obtained by dilating $\mathbb{G}_{c}$ with a circular disk area with a radius of $S$. For example, $\mathcal{M}_{\text{c}}$ is the yellow area in Fig. 5.
Unlike most MCD-based coverage path planning problems, the critical points defined here are not only on the boundary of obstacles but also on the boundary of the target regions. According to the MCD of the free space $\mathcal{W}$, a Reeb graph [21, 19] is constructed and denoted as $\mathbb{G}_{\text{w}}$. In Fig. 5, $\mathbb{G}_{\text{w}}=({\boldsymbol{C}},{\boldsymbol{E}})$ is shown as the black solid lines. The nodes of $\mathbb{G}_{\text{w}}$ are the critical points ${\boldsymbol{C}}=\{C_{i}\}$, and the edges ${\boldsymbol{E}}=\{E_{i}\}$ connect the neighboring critical points. The edge $E_{i}$ in the Reeb graph directly corresponds to the cell $\mathcal{A}_{i}$ in the free space, where $i=1,\cdots,6$ in the figure.
@@
Output : $\mathcal{P}_{R}$
1 $\mathbb{G}_{c}\leftarrow{\tt Crack\_Graph}(\mathcal{I})$, $({\boldsymbol{C}},{\boldsymbol{E}})\leftarrow{\tt MCD}(\mathcal{W}\setminus(\mathbb{G}_{c}\oplus S))$
2 $\mathbb{G}_{\text{w}}\leftarrow{\tt Reeb\_graph}({\boldsymbol{C}},{\boldsymbol{E}})$, $\mathcal{P}_{\text{wc}}\leftarrow{\tt GCC}(\mathbb{G}_{\text{w}}\cup\mathbb{G}_{c})$
3 $\textit{conn}\leftarrow{\tt cell\_connect}(\mathcal{P}_{\text{wc}},{\boldsymbol{E}})$
4 $\mathcal{P}_{R}\leftarrow{\tt complete\_coverage}(\mathcal{P}_{\text{wc}},{\boldsymbol{E}},\textit{conn},S)$
@@ 3 $\textit{conn}\leftarrow{\tt cell\_connect}(\mathcal{P}_{\text{wc}},{\boldsymbol{E}})$
Algorithm 3 briefly describes the ${\tt SCC}$ planner. With $\mathbb{G}_{\text{w}}$ and $\mathbb{G}_{c}$, we search for the shortest cost route covering all the graph edges at least once; that is, we connect edges in both graphs to form an Euler tour with the least cost. In the case of an unspecified ending position, all of the graph nodes except for the initial and ending vertices must have an even number of connected edges (i.e., an even degree); otherwise, the robot could get stuck at vertices with odd degrees. Therefore, the degree of vertices is maintained even by adding edges. For example, as shown in Fig. 5, $C_{1}$-$C_{6}$, $N_{1}$, and $N_{3}$-$N_{5}$ are the vertices with odd degrees, and the red dashed edges connecting {$C_{2}$,$N_{1}$}, {$C_{3}$,$N_{3}$}, {$C_{5}$,$N_{4}$}, and {$C_{4}$,$N_{5}$} are added to the graph to form the shortest route. The resulting Euler tour, denoted as $\mathcal{P}_{\text{wc}}$, provides an order of edges that the robot should visit (line 2). The connecting distance between each adjacent cell in the Euler tour is minimized in the function ${\tt cell\_connect}$. The variable conn represents the connections of adjacent cells in the Euler tour (line 3). Finally, complete coverage is performed by following the sequence of edges in the Euler tour, resulting in the robot path $\mathcal{P}_{R}$ (line 4).
@@ IV-B ${\tt oSCC}$ Planner
When target information is unknown, robot $\mathbf{R}$ needs to detect all targets and the critical points online in $\mathcal{W}$. We first combine graphs $\mathbb{G}_{\text{c}}$ and $\mathbb{G}_{\text{w}}$ using the following lemma and proposition with proofs given in Appendices A and B, respectively.
@@
Each critical point generated by $\mathcal{M}_{\text{c}}$ corresponds to one node of $\mathbb{G}_{\text{c}}$.
@@
Optimality is preserved with the choice of connecting critical points of $\mathcal{M}_{\text{c}}$ to corresponding nodes in $\mathbb{G}_{\text{c}}$.
@@
Figure 6: An illustrative planning example by the ${\tt oSCC}$ algorithm. (a)-(c) The Reeb graphs as additional cracks are detected during the loop cycles of the ${\tt oSCC}$ algorithm. All covered cracks are marked by thick, dotted black curves. The robot starts from the top-left corner. Each Reeb graph is updated when the robot is at point $A$ (highlighted in a circle). Those Reeb graphs are constructed after removing already-covered areas (yellow shaded areas). The red dots ($C_{i}$) denote the nodes of the Reeb graph. The solid curves ($E_{i}$) are the edges. The doubled edges are indicated by the thin, dashed black curves ($E_{i}^{\prime}$). The bold boundaries are the cell boundaries. The orange-dash lines are the robot paths. The blue-dotted dashed lines are the connections between two nodes. (d) The final path using the ${\tt oSCC}$ algorithm. The final MCD of the free space is plotted in the thick orange curves. The final trajectory is shown by the black dashed lines. The robot starts from the top-left corner and ends at the bottom left, marked by a triangle and an arrow, respectively.
@@
Output : $\mathcal{P}_{R}$
1 $\mathcal{M}_{\text{cov}}\leftarrow\O$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$
while $\mathcal{M}_{\text{cov}}\neq\mathcal{W}$ do
@@ 1 $\mathcal{M}_{\text{cov}}\leftarrow\O$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$
      3 $\mathbb{G}_{\text{w}}\leftarrow{\tt Reeb\_graph}({\boldsymbol{C}},{{\boldsymbol{E}}})$, ${\boldsymbol{\Pi}}_{w}\leftarrow{\tt Reeb\_seq}(\mathbb{G}_{\text{w}})$
      4 $\textit{conn}\leftarrow{\tt cell\_connection}({\boldsymbol{\Pi}}_{w},{{\boldsymbol{E}}})$
      5 $\mathcal{P}_{R}\leftarrow{\tt complete\_coverage}({\boldsymbol{\Pi}}_{w},{{\boldsymbol{E}}},\textit{conn},S)$
       while $true$ do